OpenStudy (anonymous):

Which point lies on the line described by the equation below? y - 3 = 2(x + 5) A. (3, 17) B. (-2, 13) C. (-1, 9) D. (3, -3) E. (4, 21) F. (4, 18) (well its not (-5, 3) so idk)

OpenStudy (anonymous):

So you start off by changing the problem to slope format.. y=mx+b

OpenStudy (anonymous):

y - 3 = 2(x + 5) y - 3 = 2x + 10 y = 2x + 13 like this?

OpenStudy (anonymous):

Yup. So then you have to graph it out. Then find out which point lies on the line.

OpenStudy (anonymous):

I actually figured it out (x Its E? :o because 21 = 2(4) + 13

OpenStudy (anonymous):

I got E. as my answer. E. (4, 21)

OpenStudy (anonymous):

You got it correct :D !! There you go.

OpenStudy (anonymous):

okay good, i thought i was just fooling myself xD aha! thankyou for your help tho!(:

OpenStudy (anonymous):

You are very welcome. If you would need help in the future, Fan me now to know if I'm available/online to assist you. -- The Wonderful Adele